# Diffusion Policy Policy Optimization

### v1

@@ 5.1 Comparison to diffusion-based RL algorithms
[…]
Overall, DPPO performs consistently, exhibits great training stability, and enjoys strong fine-tuning success rates across tasks. In the Gym tasks (Figure 5, top row), IDQL and DAWR exhibit competitive performance, while the other methods perform worse and train less stably. DPPO is the strongest performer in the Robomimic tasks (Figure 5, bottom row), especially in the challenging Transport tasks. Surprisingly, DRWR is a strong baseline in {Lift, Can, Square} but underperforms in Transport, while all other baselines fare worse still. We postulate that the other baselines, all performing off-policy Q-function-based updates, can suffer from training instability in sparse-reward Robomimic tasks especially with large action chunk sizes (see further studies in Section B.1).
Figure 5: Comparing to other diffusion-based RL algorithms. Top row: Gym tasks [8]. Bottow row: Robomimic tasks [48] with state observation. Results are averaged over five seeds in Gym tasks and three seeds in Robomimic tasks. We opt to not show the erratic error bar with QSM in Hopper-v2. DPPO curves are slightly thicker for better visualization.
@@ 5.2 Comparison to other policy parameterizations
We compare DPPO with popular RL policy parameterizations: unimodal Gaussian with diagonal covariance [76] and Gaussian Mixture Model (GMM) [5], using either MLPs or Transformers [79], and also fine-tuned with the PPO objective. We compare these to DPPO-MLP and DPPO-UNet, which use either MLP or UNet as the network backbone. We evaluate on the four tasks from Robomimic (Lift, Can, Square, Transport) with both state and pixel input. With state input, DPPO pre-trains with 20 denoising steps and then fine-tunes the last 10. With pixel input, DPPO pre-trains with 100 denoising steps and then fine-tunes 5 DDIM steps.
@@
Fig. 6 shows the results in the more challenging Square and Transport — we defer the results in Lift and Can to Fig. A8. With state input, we find that DPPO outperforms Gaussian and GMM policies, with faster convergence to $\sim$100success rate in Lift and Can, and greater final performance on Square and the challenging Transport, where it reaches $>90\%$. UNet and MLP variants perform similarly, with the latter training somewhat more rapidly. With pixel inputs, we use a Vision-Transformer-based (ViT) image encoder introduced in Hu et al. [31] and an MLP head and compare the resulting variants DPPO-ViT-MLP and Gaussian-ViT-MLP (we omit GMM due to poor performance in state-based training). While the two are comparable on Lift and Can, DPPO trains more quickly and to higher accuracy on Square, and drastically outperforms on Transport, whereas Gaussian does not improve from its 0% pre-trained success rate. To our knowledge, DPPO is the first RL algorithm to solve Transport from either state or pixel input to high (>50%) success rates.
@@ 5.3 Evaluation on Furniture-Bench, and sim-to-real transfer
Note that we are only using 50 human demonstrations for pre-training; we expect DPPO can leverage additional human data (better state space coverage) to further improve in Med, which is corroborated by ablation studies in Section B.2 examining the effect of pre-training data on DPPO.
@@
We evaluate DPPO and Gaussian policies trained in the simulated One-leg task on physical hardware zero-shot (i.e., no real data fine-tuning / co-training) over 20 trials. Please see additional simulation training and hardware details in Section D.7. Fig. 7 (bottom row) shows simulated and hardware success rates after pre-training and fine-tuning. Notably, DPPO improves the real-world performance to 80% (16 out of 20 trials). Though the Gaussian policy achieves a high success rate in simulation after fine-tuning (88%), it fails entirely on hardware (0%). Supplemental video suggests it exhibits volatile and jittery behavior. For stronger comparison, we also fine-tune the Gaussian policy with an auxiliary behavior-cloning loss [78] such that the fine-tuned policy is encouraged to stay close to the base policy. However, this limits fine-tuning and only leads to a 53% success rate in simulation and 50% in reality.
Qualitatively, we find fine-tuned policies to be more robust and exhibit more corrective behaviors than pre-trained-only policies, especially during the insertion stage of the task. Fig. 8 shows frames from representative rollouts on hardware. Overall, these results demonstrate the strong sim-to-real capabilities of DPPO; Section 6 provides a conjectural mechanism for why this may be the case.
@@ 5.4 Summary of ablation findings
We conduct extensive ablation studies in Section B.1. Our main findings include: (1) for challenging tasks, using a value estimator which depends on environment state but is independent of denoised action is crucial for performance; we conjecture that this is related to the high stochasticity of Diffusion Policy; (2) there is a sweet spot for clipping the denoising noise level for DPPO exploration, trading off between too little exploration and too much action noise; (3) DPPO is resilient to fine-tuning fewer-than-$K$ denoising steps, yielding improved runtime and comparable performance; (4) DPPO yields improvements over Gaussian-MLP baselines for varying levels of expert demonstration data, and achieves comparable final performance and sample efficiency when training from scratch in Gym environments.
@@ 6 Understanding the performance of DPPO
We study the factors contributing to DPPO’s improvements in performance over the popular Gaussian and GMM methods introduced in Section 5.2. Our findings highlight three major contributing factors:
@@
We use the Avoid environment from the D3IL benchmark introduced in Jia et al. [36], where a robot arm needs to reach the other side of the table while avoiding an array of obstacles (Fig. 9, top-left). The action space is the 2D target location of the end-effector. D3IL provides a set of expert demonstrations that covers different possible paths to the goal line — we consider three subsets of the demonstrations, M1, M2, and M3 in Fig. 9, each with two distinct modes. We choose such relatively simple setups with only two modes in each setting such that Gaussian (with exploration noise) [^5] Without noise, Gaussian policy is fully deterministic and cannot capture the two modes. and GMM can fit the expert data distribution reasonably well, allowing fair comparisons in fine-tuning.
We pre-train MLP-based Diffusion, Gaussian, and GMM policies (action chunk size $T_{a}=4$ unless noted) with the demonstrations. For fine-tuning, we assign (sparse) reward when the robot reaches the goal line from the topmost mode. Gaussian and GMM policies are also fine-tuned with the PPO objective.
Figure 9: (Left) We use the Avoid environment from D3IL benchmark [36] to visualize the DPPO’s exploration tendencies. We design the task of always reaching the green goal line from the topmost mode. (Right) Structured exploration. We show the sampled trajectories at the beginning of fine-tuning for DPPO, Gaussian, and GMM after pre-training on three sets of expert demonstrations, M1, M2, and M3.
@@
Fig. 9 (right) shows the sampled trajectories from DPPO, Gaussian, and GMM at the first iteration of fine-tuning. DPPO explores in wide coverage around the expert data manifold, whereas Gaussian generates less structured exploration noise (especially in M2) and GMM exhibits narrower coverage. Moreover, the combination of diffusion parameterization with the denoising of action chunks means that policy stochasticity in DPPO is structured in both action dimension and time horizon. Quantitatively, Fig. A7 in Appendix shows DPPO achieves greater fine-tuning efficiency than Gaussian and GMM in all settings if a sufficient number of denoising steps is fine-tuned, consistent with the findings in Section 5.2 and Section 5.4.
It is possible, however, that the on-manifold exploration we observe with DPPO hinders fine-tuning when aggressive, unstructured exploration is desired. We conjecture this is the case in the Lamp environment with Low randomness, in which DPPO slightly underperforms the Gaussian baseline (Fig. 7). Moreover, we do not observe that DPPO is substantially better (nor is it any worse) than Gaussian in exploration from-scratch (see Appendix B). Thus, we anticipate that this structured, on-manifold exploration confers the greatest benefit when pre-training provides sufficient coverage of relevant success modes. In particular, we believe that this makes DPPO uniquely suited for fine-tuning large Diffusion Policy pre-trained on multiple tasks [83], as it may exhibit sufficient mode coverage given training data diversity.
@@
In Fig. 10 (left), we run fine-tuning after pre-training with M2 and attempt to de-stabilize fine-tuning by gradually adding noise to the action during the fine-tuning process (see Section D.8 for details). We find that Gaussian and GMM’s performance both collapse, while with DPPO, the performance is robust to the noise if at least four denoising steps are used. This property also allows DPPO to apply significant noise to the sampled actions, simulating an imperfect low-level controller to facilitate sim-to-real transfer in Section 5.3. In Fig. 10 (right), we also find DPPO enjoys greater training stability when fine-tuning long action chunks, e.g., up to $T_{a}=16$, while Gaussian and GMM can fail to improve at all.
@@
DPPO also generates final policies robust to perturbations in dynamics and the initial state distribution. In Fig. 12, we again add noise to the actions sampled from the fine-tuned policy (no noise applied during training) and find that the DPPO policy exhibits strong robustness to the noise compared to the Gaussian policy. The DPPO policy also converges to the (near-)optimal path from a larger distribution of initial states. This finding echoes theoretical guarantees that Diffusion Policy, capable of representing complex multi-modal data distribution, can effectively deconvolve noise from noisy states [7], a property used in Chen et al. [11] to stabilize long-horizon video generation.
Figure 12: Policy robustness. Green dot / box indicates the initial state region.
@@ 7 Conclusion and Future Work
We believe DPPO will become an important component in the pre-training-plus-fine-tuning pipeline for training general-purpose real-world robotic policies. To this end, we hope in future work to further showcase the promise of DPPO for simulation-to-real transfer [41, 62, 14] in which we fine-tune a vision-based policy that has been pre-trained on a variety of diverse tasks. We expect this pre-training to provide a large and diverse expert data manifold, of which, as we have shown in Section 6, DPPO is well-suited to take advantage of better exploration during fine-tuning. We are also excited to understand how DPPO can fit together with other decision-making tools, such as model-based planning [35] and decision-making aided by video prediction [11]. Finally, as noted in the introduction, we eagerly anticipate applications of DPPO in domains beyond robotics, where diffusion models have shown promise for combinatorial search and sequence modeling [58, 45].

### v3

@@ 5.1 Comparison to diffusion-based RL algorithms
We compare DPPO to an extensive list of RL methods for fine-tuning diffusion-based policies. Baseline names are color-coordinated with their plot colors. Two methods, DRWR, DAWR, are our own, novel baselines that are based on reward-weighted regression (RWR, Peters and Schaal [70]) and advantage-weighted regression (AWR, Peng et al. [67]). The remaining methods, DIPO [103], IDQL [35], DQL [100], and QSM [74], are existing in the literature. Among them, QSM and DIPO are proposed specifically for training-from-scratch, while IDQL and DQL can be more suitable to online fine-tuning. We evaluate on the three OpenAI Gym tasks and the four Robomimic tasks with state input; detailed descriptions of all baselines and training details are in Section E.3.
Overall, DPPO performs consistently, exhibits great training stability, and enjoys strong fine-tuning performance across tasks. In the Gym tasks (Figure 5, top row), IDQL and DIPO exhibit competitive performance, while the other methods often perform worse and train less stably. Baselines also tend to exhibit performance drop at the beginning of fine-tuning, mostly likely due to misestimating Q values of the pre-trained policy. DPPO is the strongest performer in the Robomimic tasks (Figure 5, bottom row), especially in the challenging Transport tasks. IDQL and, surprisingly, DRWR, are strong baselines in {Lift, Can, Square} but underperforms in Transport, while all other baselines fare worse still. We postulate that the other baselines, all performing off-policy updates and propagating gradients from the imperfect Q function to the actor, suffers from even greater training instability in sparse-reward Robomimic tasks given the continuous action space plus large action chunk sizes (see furtuer studies in Section C.2).
Figure 5: Comparing to other diffusion-based RL algorithms. Top row: Gym tasks [11] averaged over five seeds. Bottow row: Robomimic tasks [60], averaged over three seeds, with state observation. DPPO curves are slightly thicker for better visibility.
5.2 Comparison to other demo-augmented RL algorithms
We compare DPPO with recently proposed RL methods for training robot policies (not necessarily diffusion-based) leveraging offline data, including RLPD [7], Cal-QL [62], and IBRL [40]. These methods add expert data in the replay buffer and performs off-policy updates (IBRL and Cal-QL also pretrain with behavior cloning and offline RL objectives, respectively), which significantly improves sample efficiency vs. DPPO in HalfCheetah-v2 (Fig. 6, top left).
Among the Franka-Kitchen settings (Fig. 6, top right), we find RLPD and IBRL fail to learn well especially with noisier demonstrations from Kitchen-Partial-v0 and Kitchen-Mixed-v0. Cal-QL achieves competitive performance but DPPO still achieves overall the best performance especially with Kitchen-Complete-v0. We note that DPPO, not using any expert data during fine-tuning, can be sensitive to the pre-training performance; we find the incomplete demonstrations in Kitchen-Partial-v0 and Kitchen-Mixed-v0 cause challenge in fully modeling the multi-modality of the data even with diffusion parameterization and prevent DPPO from achieving (near-)perfect fine-tuning performance.
Nonetheless, we believe the expert demonstrations from Robomimic are most reflective of pre-training plus fine-tuning in robotics as all demonstrations complete the task despite the varying quality. Fig. 6 bottom row shows the performance of DPPO and baselines using either cleaner PH or noisier MH data in Can and Square; DPPO exhibits much stronger final performance. RLPD and Cal-QL fail to learn at all and IBRL saturates at lower success levels. DPPO also runs significantly faster in wall-clock time than the baselines as it leverages sampling from highly parallelized environments [^6] Off-policy methods (baselines) usually cannot fully leverage parallelized sampling as the policy is updated less often (e.g., 50 updates per 50 samples instead of 1 update per 1 sample) and the performance can be affected.; thus we cap the number of samples at 1e6 for the baselines in Robomimic, also since their performance saturates.
Figure 6: Comparing to other demo-augmented RL algorithms. Top left: HalfCheetah-v2 task [11] averaged over five seeds. Top right: Franka-Kitchen task [32] averaged over three seeds. Bottom row: Robomimic tasks [60] averaged over three seeds with state observation, using either Proficient-Human (PH) or Multi-Human (MH) data.
@@ 5.3 Comparison to other policy parameterizations
We compare DPPO with popular RL policy parameterizations: unimodal Gaussian with diagonal covariance [91] and Gaussian Mixture Model (GMM [8]), using either MLPs or Transformers [94], and also fine-tuned with the PPO objective. We compare these to DPPO-MLP and DPPO-UNet, which use either MLP or UNet as the network backbone. We evaluate on the four tasks from Robomimic (Lift, Can, Square, Transport) with both state and pixel input. With state input, DPPO pre-trains with 20 denoising steps and then fine-tunes the last 10. With pixel input, DPPO pre-trains with 100 denoising steps and then fine-tunes 5 DDIM steps.
@@
Fig. 7 display results for the more challenging Square and Transport — we defer the results in Lift and Can to Fig. A9. With state input, DPPO outperforms Gaussian and GMM policies, with faster convergence to $\sim$100success rate in Lift and Can, and greater final performance on Square and the challenging Transport, where it reaches $>90\%$. UNet and MLP variants perform similarly, with the latter training somewhat more rapidly. With pixel inputs, we use a Vision-Transformer-based (ViT) image encoder introduced in Hu et al. [40] and an MLP head and compare the resulting variants DPPO-ViT-MLP and Gaussian-ViT-MLP (we omit GMM due to poor performance in state-based training). While the two are comparable on Lift and Can, DPPO trains more quickly and to higher accuracy on Square, and drastically outperforms on Transport, whereas Gaussian does not improve from its 0% pre-trained success rate. To our knowledge, DPPO is the first RL algorithm to solve Transport from either state or pixel input to high (>50%) success rates.
@@ 5.4 Evaluation on Furniture-Bench, and sim-to-real transfer
Note that we are only using 50 human demonstrations for pre-training; we expect DPPO can leverage additional human data (better state space coverage) to further improve in Med, which is corroborated by ablation studies in Section C.3 examining the effect of pre-training data on DPPO.
@@
We evaluate DPPO and Gaussian policies trained in the simulated One-leg task on physical hardware zero-shot (i.e., no real data fine-tuning / co-training) over 20 trials. Both policies run at 10Hz. Please see additional simulation training and hardware details in Section E.8. Fig. 8 (bottom row) shows simulated and hardware success rates after pre-training and fine-tuning. Notably, DPPO improves the real-world performance to 80% (16 out of 20 trials). Though the Gaussian policy achieves a high success rate in simulation after fine-tuning (88%), it fails entirely on hardware (0%). Supplemental video suggests it exhibits volatile and jittery behavior. For stronger comparison, we also fine-tune the Gaussian policy with an auxiliary behavior-cloning loss [93] such that the fine-tuned policy is encouraged to stay close to the base policy. However, this limits fine-tuning and only leads to 53% success rate in simulation and 50% in reality.
Qualitatively, we find fine-tuned policies to be more robust and exhibit more corrective behaviors than pre-trained-only policies, especially during the insertion stage of the task; such behaviors are visualized in Fig. 2 and Fig. 9 shows representative rollouts on hardware. Overall, these results demonstrate the strong sim-to-real capabilities of DPPO; Section 6 provides a conjectural mechanism for why this may be the case.
@@ 5.5 Summary of ablation findings
We conduct extensive ablation studies in Section C.2. Our main findings include: (1) for challenging tasks, using a value estimator which depends on environment state but is independent of denoised action is crucial for performance; we conjecture that this is related to the high stochasticity of Diffusion Policy; (2) there is a sweet spot for clipping the denoising noise level for DPPO exploration, trading off between too little exploration and too much action noise; (3) DPPO is resilient to fine-tuning fewer-than-$K$ denoising steps, yielding improved runtime and comparable performance; (4) DPPO yields improvements over Gaussian-MLP baselines for varying levels of expert demonstration data, and achieves comparable final performance and sample efficiency when training from scratch in Gym environments.
@@ 6 Understanding the performance of DPPO
The improvement of DPPO over popular Gaussian and GMM methods in Section 5.3 comes as a surprise initially as DPPO solves a much longer Diffusion Policy MDP (Section 4) than the original environment MDP that other methods solve. This leads us to study the factors contributing to DPPO’s improvements in performance. Our findings highlight three major contributing factors:
@@
We use the Avoid environment from D3IL benchmark [46], where a robot arm needs to reach the other side of the table while avoiding an array of obstacles (Fig. 10, top-left). The action space is the 2D target location of the end-effector. D3IL provides a set of expert demonstrations that covers different possible paths to the goal line — we consider three subsets of the demonstrations, M1, M2, and M3 in Fig. 10, each with two distinct modes. We choose such relatively simple setups with only two modes in each setting such that Gaussian (with exploration noise) [^7] Without noise, Gaussian policy is fully deterministic and cannot capture the two modes. and GMM can fit the expert data distribution reasonably well, allowing fair comparisons in fine-tuning.
We pre-train MLP-based Diffusion, Gaussian, and GMM policies (action chunk size $T_{a}=4$ unless noted) with the demonstrations. For fine-tuning, we assign sparse reward when the robot reaches the goal line from the topmost mode. Gaussian and GMM policies are also fine-tuned with the PPO objective.
Figure 10: (Left) We use the Avoid environment from D3IL benchmark [46] to visualize the DPPO’s exploration tendencies. We design the task of always reaching the green goal line from the topmost mode. (Right) Structured exploration. We show sampled trajectories at the first iteration of fine-tuning for DPPO, Gaussian, and GMM after pre-training on three sets of expert demonstrations, M1, M2, and M3.
@@
Fig. 10 (right) shows the sampled trajectories (with exploration noise) from DPPO, Gaussian, and GMM during the first iteration of fine-tuning. DPPO explores in wide coverage around the expert data manifold, whereas Gaussian generates less structured exploration noise (especially in M2) and GMM exhibits narrower coverage. Unlike Gaussian policy that adds noise only to the final sampled action, diffusion adds multiple rounds of noise through denoising. Each denoising step expands the coverage with new noise while also pushing the newly denoised action towards the expert data manifold [69]. Moreover, the combination of diffusion parameterization with the denoising of action chunks means that policy stochasticity in DPPO is structured in both action dimension and time horizon. Quantitatively, Fig. A8 in Appendix shows DPPO achieves greater fine-tuning efficiency than Gaussian and GMM in all settings if a sufficient number of denoising steps is fine-tuned, consistent with the findings in Section 5.3 and Section 5.5.
It is possible, however, that the on-manifold exploration we observe with DPPO hinders fine-tuning when aggressive, unstructured exploration is desired. We conjecture this is the case in the Lamp environment with Low randomness, in which DPPO slightly underperforms the Gaussian baseline (Fig. 8). Moreover, we do not observe that DPPO is substantially better (nor is it any worse) than Gaussian in exploration from-scratch (see Appendix C). Thus, we anticipate that this structured, on-manifold exploration confers the greatest benefit when pre-training provides sufficient coverage of relevant success modes. In particular, we believe that this makes DPPO uniquely suited for fine-tuning large Diffusion Policy pre-trained on multiple tasks [99], as it may exhibit sufficient mode coverage given training data diversity.
@@
In Fig. 11 (left), we run fine-tuning after pre-training with M2 and attempt to de-stabilize fine-tuning by gradually adding noise to the action during the fine-tuning process (see Section E.9 for details). We find that Gaussian and GMM’s performance both collapse, while with DPPO, the performance is robust to the noise if at least four denoising steps are used. This property also allows DPPO to apply significant noise to the sampled actions, simulating an imperfect low-level controller to facilitate sim-to-real transfer in Section 5.4. In Fig. 11 (right), we also find DPPO enjoys greater training stability when fine-tuning long action chunks, e.g., up to $T_{a}=16$, while Gaussian and GMM can fail to improve at all.
@@
DPPO also generates final policies robust to perturbations in dynamics and the initial state distribution. In Fig. 13, we again add noise to the actions sampled from the fine-tuned policy (no noise applied during training) and find that DPPO policy exhibits strong robustness to the noise compared to the Gaussian policy. DPPO policy also converges to the (near-)optimal path from a larger distribution of initial states. This finding echoes theoretical guarantees that Diffusion Policy, capable of representing complex multi-modal data distribution, can effectively deconvolve noise from noisy states [10], a property used in Chen et al. [15] to stabilize long-horizon video generation.
Figure 13: Policy robustness after fine-tuning. Green dot / box indicates the initial state region.
@@ 7 Conclusion and Future Work
We believe DPPO will become an important component in the pre-training-plus-fine-tuning pipeline for training general-purpose real-world robotic policies. To this end, we hope in future work to further showcase the promise of DPPO for simulation-to-real transfer [18, 52, 77, 19] in which we fine-tune a vision-based policy that has been pre-trained on a variety of diverse tasks. We expect this pre-training to provide a large and diverse expert data manifold, of which, as we have shown in Section 6, DPPO is well-suited to take advantage for better exploration during fine-tuning. We are also excited to understand how DPPO can fit together with other decision-making tools such as model-based planning [45] and decision-making aided by video prediction [15]; these tools may help address the main limitation of DPPO — its lower sample efficiency than off-policy methods — and unlocking performing practical RL in physical hardware or generative simulation (e.g., video-based [13, 104]). Finally, as noted in the introduction, we eagerly anticipate applications of DPPO in domains beyond robotics, where diffusion models have shown promise for combinatorial search and sequence modeling [73, 56].
@@ Appendix A Extended Related Work
A.1 RL training of robot policies with offline data
Here, we discuss related work in training robot policies using RL augmented with offline data to help RL better explore online in sparse reward settings.
One simple form is to use offline data to pre-train the policy, typically using behavior cloning, and then fine-tune the policy online. This is the approach that DPPO takes. Often, a regularization loss is applied to constrain the fine-tuned policy to stay close to the base policy, leading to natural fine-tuned behavior and often better learning [76, 107, 93]. DPPO does not apply regularization at fine-tuning as we find the on-manifold exploration helps DPPO maintain natural behavior after fine-tuning Section 5.4. Another popular approach is to learn a residual policy with RL on top of the frozen base policy [3, 34]. A closer work to ours is from Ankile et al. [6], which trains a one-step residual non-diffusion policy with on-policy RL on top of a pre-trained chunked diffusion policy. This approach has the benefit of being fully closed-loop but lacks the structured on-manifold exploration of DPPO. Another hybrid approach is from Hu et al. [40], which uses pre-trained and fine-tuned policies to sample online experiences.
Another popular line of work, instead of training a base policy using offline data, directly adds the data in the replay buffer for online, off-policy learning in a single stage [37, 95, 61]. One recent approach from Ball et al. [7], RLPD, further improves sample efficiency from previous off-policy methods incorporating, e.g., critic ensembling. Luo et al. [57] demonstrate RLPD solving real-world manipulation tasks (although generally less challenging than ones solved by DPPO).
Other approaches, including Cal-QL, build on offline RL to learn from offline data and then switch to online RL while still sampling from offline data [62, 35, 102]. Often the distributional mismatch between offline data and online policy needs to be addressed: Cal-QL proposes calibrated conservative Q-learning that learns a offline Q function that lower bounds the true value of the learned policy; Lei et al. [51] propose ensemble behavior cloning during pre-training to promote policy diversity.
A.2 Diffusion-based RL methods